An object initially at rest accelerates uniformly at 5 m/sec west decelerates uniformly at 4.0 m/sec^2 for 2.0 sec. What is the displacement of this object?
1
Expert's answer
2010-10-22T02:50:23-0400
Let's write the equations of the motion:
h = h0 +v0 t +at2/2;
For our case:
h = 5 * 2 - 4*22/2 = 2 m.
The displacement is 2 m.